OpenStudy (anonymous):

It costs $2.80 to make a sandwich at the local deli shop. To make a profit, the deli sells it at a price that is 170% of the cost. The sandwich sells for $___.

OpenStudy (anonymous):

2.80 * 1.7 = $4.76

OpenStudy (kash_thesmartguy):

Let's see:\[2.80*1.7=\]Add the number to 2.80 to get your answer

OpenStudy (kash_thesmartguy):

SO, add 4.76 to 2.80 to get your answer!

OpenStudy (kash_thesmartguy):

\[2.80+4.76=YourAnswer\]

OpenStudy (anonymous):

thanks

OpenStudy (kash_thesmartguy):

:)